given a function 4x-4 what is y when x is 1, 2, and 3



1
Expert's answer
2021-08-23T17:03:22-0400

Given that the function is 4x-4.

Therefore, y=f(x)=4x-4

Now, when x=1, we get y=4(1)-4=4-4=0

When x=2, we get y=4(2)-4=8-4=4

When x=3, we get y=4(3)-4=12-4=8